what is x in the question solve for x in the problem 3/7(42 x+28)=-5/8(16 x+48)+4
Answer by Cromlix(4381)

Hi there,
3/7(42x+28)=-5/8(16x+48)+4
Divide brackets by 7 and 8
3(6x + 4) = -5(2x + 6) + 4
18x + 12 = -10x - 30 + 4
18x + 10x = -30 + 4 - 12
28x = -38
x = -38/28
x = -19/14
x = -1 5/14
Hope this helps :-)